OpenStudy (anonymous):

a cash register contains 18 quarters and dimes, find the total of the coins in the terms of quarters, q

OpenStudy (sandra):

ok, so they're asking you for an equation that represent the relationship between quarters and dimes

OpenStudy (sandra):

so if there are 18 total coins, then we know that #quarters + #dimes = 18

OpenStudy (sandra):

so if d+q=18, then d+q-d = 18-d, and q=18-d

OpenStudy (sandra):

that is, the number of quarters must be 18 minus however many dimes there are, since you know there are 18 coins and only those two types